OpenStudy (rawfeeds-):

Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y –3 7 –2 5 Line 2 x y 0 5 2 1 Line 1: y = x + Line 2: y = x +

OpenStudy (rawfeeds-):

@ijlal @sweetburger

OpenStudy (ijlal):

use the method that i used in your last posted question find the slope then put in y-y1=m(x-x1)

OpenStudy (ijlal):

m=(y2-y1)/(x2-x1)

OpenStudy (jdcap):

\[y_1=1-2x\\y_2=5-2x\]

OpenStudy (rawfeeds-):

ok

OpenStudy (rawfeeds-):

is that th anwser

OpenStudy (rawfeeds-):

hello @ijlal

OpenStudy (rawfeeds-):

omg

OpenStudy (ijlal):

a moment

OpenStudy (rawfeeds-):

@ijlal

OpenStudy (rawfeeds-):

@iGreenHacked

OpenStudy (ijlal):

yep those are the lines that @JDcap stated

OpenStudy (ijlal):